# for the Hadron Spectrum Collaboration

### v1

@@ V.2 Poles in coupled-channel ${DD^{\ast}}-{D^{\ast}{D}^{\ast}}$ S-wave scattering
In coupled-channel ${DD^{\ast}}$-${D^{\ast}{D}^{\ast}}$ scattering, the complex $s$-plane is divided into 4 sheets due to the branch cut present at each threshold. The sheets are labelled by the sign of the imaginary part of the cm-frame momentum for each hadron channel using the notation $(\mathrm{sign}[\mathrm{Im}(k_{{DD^{\ast}}}]),\mathrm{sign}[\mathrm{Im}(k_{{D^{\ast}{D}^{\ast}}})])$. The physical sheet, where physical scattering occurs, is defined by $\mathrm{Im}(k_{{DD^{\ast}}}),\mathrm{Im}(k_{{D^{\ast}{D}^{\ast}}})>0$. We refer to this as sheet I, or equivalently the (+,+)(+,+)( + , + ) sheet. Causality forbids that poles off the real energy axis occur on the physical sheets. However, they are allowed to occur on the unphysical sheets II (,+)(-,+)( - , + ), III (,)(-,-)( - , - ) and IV (+,)(+,-)( + , - ) where they occur as complex-conjugate pairs in $s$. It is poles on these unphysical sheets that correspond to resonances with $\sqrt{s_{0}}=m\pm\frac{i}{2}\Gamma$, where $m$ is the mass of the resonance and $\Gamma$ its width. A pole off the real energy axis on an unphysical sheet can influence physical scattering if it is close to the physical sheet. In the case of sheets II and III, it is the lower half plane that is in close proximity, whereas for sheet IV it is the upper half. We will denote the relevant sheets using the shorthand $\mathrm{I}_{u},\mathrm{II}_{l},\mathrm{III}_{l}$ and $\mathrm{IV}_{u}$. Close to the pole, the scattering $t$-matrix is of the form of Eq. (12).
@@ V.3 Varying the ${DD^{\ast}}\rightarrow{D^{\ast}{D}^{\ast}}$ $S$-wave Amplitude
Figure 11 shows the amplitudes (top) and poles position (bottom) for each value of the coupling parameter. We observe that as the coupling parameter decreases, the near threshold cusp in the ${DD^{\ast}}$ $S$-wave amplitude becomes less pronounced, eventually disappearing when the coupling is completely turned off. This is accompanied by a rapid opening of the ${D^{\ast}{D}^{\ast}}$ $S$-wave amplitude. The behavior of the amplitude can be explained by examining the position of the sheet IV pole. As the coupling parameter is decreased, the sheet IV pole moves to the axis with purely imaginary momentum in conjunction with the disappearance of the cusp. We interpret this resonance as a virtual bound state in the limit that the two hadron scattering channels are decoupled, which correspondingly enhances the ${D^{\ast}{D}^{\ast}}$ $S$-wave amplitude upon opening. However, due to the open ${DD^{\ast}}$ channel and nonzero coupling, the pole is moved to a complex energy, manifesting as a resonance which is responsible for the cusp in the ${DD^{\ast}}$ $S$-wave amplitude.
Also examined is the behavior of the sheet II and sheet III poles as the hadron-hadron channels are decoupled. Table 6 shows the momentum of the sheet II and III poles in the ${D^{\ast}{D}^{\ast}}$ plane for the different values of the coupling parameter. When the two channels are fully decoupled, the sheet III pole becomes an exact “mirror” of the sheet II pole with opposite sign of momentum. Due to this behavior and in conjunction with its distance from the physical sheet, we do not prescribe it any physical significance.
@@ V.4 Poles in other $J^{P}$
where the quoted uncertainities are statistical. We stress that this is just a single parameterisation and so no firm conclusions should be drawn. With such a small amount of constraint, higher order amplitudes can easily introduce complex poles on the physical sheet [^4] This occurs in all amplitudes attempted with a $K$-matrix with a linear polynomial in $s$.. Given that relatively few energy levels are available, we do not investigate this further. However, it is possible this resonance pole is a feature in $J^{P}=0^{-}$ ${DD^{\ast}}\{{\>\!}^{{3}\!}{P}_{0}\}$ scattering. The width is large, more than 300 MeV in physical units.
Other more distant poles can be found in some of the other background partial waves. We do not consider any of these important as they are far away from where we constrain the scattering amplitudes and have a negligible effect on physical scattering.
@@ VI.1 Systematic Effects and Left-Cut Physics
The effects of left-cut physics arising from meson exchange in the scattering $t,u$-channels have not been accounted for in this calculation. They are also not currently included in the most general finite volume formalism applicable to coupled-channel scattering. The finite volume energy levels should include the necessary physics, but this has not been included in translating the lattice QCD energy levels into scattering information. At the masses used in this calculation, a $u$-channel cut due to $\pi$ exchange [^5] This is the equivalent of the “short nucleon cut” cut seen in $N\pi$ scattering ends at $a_{t}E_{lhc}=0.68432$, roughly 18 MeV below threshold, and only a little below the lowest energy level extracted ($a_{t}E=0.68474(46)$ in $[000]T_{1}^{+}$ on the $L/a_{s}=16$ lattice). The virtual bound state pole we have identified would lie on this cut. The effect occurs again a similar distance below ${D^{\ast}{D}^{\ast}}$ threshold and here we have many energy levels overlapping this cut from the elastic ${DD^{\ast}}$ region. Some modification is likely required to account for these cuts; at the very least, small imaginary parts must be introduced in the scattering amplitudes. However, as the amplitudes are consistent with the finite volume spectrum, it is possible that the effects of the left-cut physics are not that large.
Efforts are underway to add the necessary terms to the two-body Lüscher formalism to include these effects [64, 65]. A perhaps more general alternative appears to be to continue the three-body formalism a short way below threshold to describe this system [66].
Lippmann-Schwinger and effective field theory methods applied to the ${DD^{\ast}}$ system at larger-than-physical pion masses, where the $D^{\ast}$ is stable, can be found in Refs. [67, 68, 43]. There it is shown that the presence of the left-cut below ${DD^{\ast}}$ threshold arising from one pion exchange affects both the nature of the pole and its position. Similar complications can be expected below ${D^{\ast}{D}^{\ast}}$ threshold, and so extra work is needed to understand how the structures observed in this study are affected.
@@ VI.2 Comparison to Previous Work and Phenomenology
Other works have studied this channel. Considering only elastic scattering, Ref. [42] working at a light-quark mass corresponding to $m_{\pi}=280$ MeV, finds a virtual bound state 9.9 MeV below ${DD^{\ast}}$ threshold in $S$-wave. Using the HAL QCD method with almost-physical pions, Ref. [40] finds a virtual bound state much closer to threshold. While these results point toward a ground-state $T_{cc}$ that gets closer to threshold with decreasing light-quark mass, further work is needed to understand the structures in the complex plane before firm conclusions can be made.
@@ VII Summary
In summary, we have presented the first coupled-channel ${DD^{\ast}}$, ${D^{\ast}{D}^{\ast}}$ scattering calculation from lattice QCD in $S$-wave for isospin-0. Using 109 energy levels, we determined the $J^{P}=1^{+}$ scattering amplitude by constraining background waves in the ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ channels for 12 parameterization variations that adequately describe the finite volume spectrum. In all successful variations, we observe a pole in the $S$-wave amplitudes below ${DD^{\ast}}$ threshold corresponding to a virtual bound state at approximately $\sqrt{s}\approx 3834(31)$ MeV, commonly identified as the $T_{cc}$. A pole corresponding to a resonance $T_{cc}^{\prime}$ is also found just below ${D^{\ast}{D}^{\ast}}$ threshold at $\sqrt{s}\approx 3971(35)$ MeV with a width of $\Gamma\approx 11(13)$ MeV, which predominantly couples to the kinematically closed ${D^{\ast}{D}^{\ast}}$ channel. The presence of the nearby resonance pole below ${D^{\ast}{D}^{\ast}}$ threshold enhances the ${DD^{\ast}}$ $S$-wave amplitude, resulting in clear cusp at ${D^{\ast}{D}^{\ast}}$ threshold. As ongoing experiments are currently investigating the isospin-0 $J^{P}=1^{+}$ charm sector, it should be possible to find the experimental $T_{cc}^{\prime}$ in both ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ final states.
@@ Acknowledgements.
We thank our colleagues within the Hadron Spectrum Collaboration (www.hadspec.org), particularly Jozef Dudek, Raul Briceño and Andrew Jackura for useful discussions. TW & DJW acknowledges support from a Royal Society University Research Fellowship. TW, DJW & CET acknowledge support from the U.K. Science and Technology Facilities Council (STFC) [grant numbers ST/T000694/1, ST/X000664/1]. TW also acknowledges support from Science Foundation Ireland [grant number 21/FFP-P/10186] The software codes Chroma [71], QUDA [72, 73], QUDA-MG [74], QPhiX [75], MG_PROTO [76], QOPQDP [77, 78], and Redstar [79] were used. Some software codes used in this project were developed with support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing (SciDAC) program; also acknowledged is support from the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration. This work used the Cambridge Service for Data Driven Discovery (CSD3), part of which is operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk) on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 was funded by BEIS capital funding via STFC capital grants ST/P002307/1 and ST/R002452/1 and STFC operations grant ST/R00689X/1. Other components were provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/P020259/1). This work also used the earlier DiRAC Data Analytic system at the University of Cambridge. This equipment was funded by BIS National E-infrastructure capital grant (ST/K001590/1), STFC capital grants ST/H008861/1 and ST/H00887X/1, and STFC DiRAC Operations grant ST/K00333X/1. DiRAC is part of the National E-Infrastructure. This work also used clusters at Jefferson Laboratory under the USQCD Initiative and the LQCD ARRA project. Propagators and gauge configurations used in this project were generated using DiRAC facilities, at Jefferson Lab, and on the Wilkes GPU cluster at the University of Cambridge High Performance Computing Service, provided by Dell Inc., NVIDIA and Mellanox, and part funded by STFC with industrial sponsorship from Rolls Royce and Mitsubishi Heavy Industries. Also used was an award of computer time provided by the U.S. Department of Energy INCITE program and supported in part under an ALCC award, and resources at: the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725; the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231; the Texas Advanced Computing Center (TACC) at The University of Texas at Austin; the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation Grant No. ACI-1548562; and part of the Blue Waters sustained-petascale computing project, which is supported by the National Science Foundation (awards OCI-0725070 and ACI-1238993) and the state of Illinois. Blue Waters is a joint effort of the University of Illinois at Urbana-Champaign and its National Center for Supercomputing Applications.
Data access statement
@@ Appendix A Subduction Tables
In this Appendix, we list the pseudoscalar-vector and identical vector-vector partial waves that contribute to the finite volume irreducible representations. Note that for identical vector-vector scattering, certain partial waves are not allowed by Bose symmetry (see Table 2) and these are ommitted from the table.
@@ Appendix B Operator Tables
${D^{\ast}{D}^{\ast}}$ PW $\gamma^{(0)}_{{D^{\ast}{D}^{\ast}}\{{\>\!}^{{2S+1}\!}{\ell}_{J}\}\rightarrow{D^{\ast}{D}^{\ast}}\{{\>\!}^{{2S+1}\!}{\ell}_{J}\}}$ $\chi^{2}/N_{\mathrm{dof}}$

### v2

@@ V.2 Poles in coupled-channel ${DD^{\ast}}-{D^{\ast}{D}^{\ast}}$ S-wave scattering
In coupled-channel ${DD^{\ast}}$-${D^{\ast}{D}^{\ast}}$ scattering, the complex $s$-plane is divided into 4 sheets due to the branch cut present at each threshold. The sheets are labelled by the sign of the imaginary part of the cm-frame momentum for each hadron channel using the notation $(\mathrm{sign}[\mathrm{Im}(k_{{DD^{\ast}}}]),\mathrm{sign}[\mathrm{Im}(k_{{D^{\ast}{D}^{\ast}}})])$. The physical sheet, where physical scattering occurs, is defined by $\mathrm{Im}(k_{{DD^{\ast}}}),\mathrm{Im}(k_{{D^{\ast}{D}^{\ast}}})>0$. We refer to this as sheet I, or equivalently the (+,+)(+,+)( + , + ) sheet. Causality forbids that poles off the real energy axis occur on the physical sheet. However, they are allowed to occur on the unphysical sheets II (,+)(-,+)( - , + ), III (,)(-,-)( - , - ) and IV (+,)(+,-)( + , - ) where they occur as complex-conjugate pairs in $s$. It is poles on these unphysical sheets that correspond to resonances with $\sqrt{s_{0}}=m\pm\frac{i}{2}\Gamma$, where $m$ is the mass of the resonance and $\Gamma$ its width. A pole off the real energy axis on an unphysical sheet can influence physical scattering if it is close to the physical sheet. In the case of sheets II and III, it is the lower half plane that is in close proximity, whereas for sheet IV it is the upper half. We will denote the relevant sheets using the shorthand $\mathrm{I}_{u},\mathrm{II}_{l},\mathrm{III}_{l}$ and $\mathrm{IV}_{u}$. Close to the pole, the scattering $t$-matrix is of the form of Eq. (12).
@@ V.3 Varying the ${DD^{\ast}}\rightarrow{D^{\ast}{D}^{\ast}}$ $S$-wave Amplitude
Figure 11 shows the amplitudes (top) and pole position (bottom) for each value of the coupling parameter. We observe that as the coupling parameter decreases, the near threshold cusp in the ${DD^{\ast}}$ $S$-wave amplitude becomes less pronounced, eventually disappearing when the coupling is completely turned off. This is accompanied by a rapid opening of the ${D^{\ast}{D}^{\ast}}$ $S$-wave amplitude. The behavior of the amplitude can be explained by examining the position of the sheet IV pole. As the coupling parameter is decreased, the sheet IV pole moves to the axis with purely imaginary momentum in conjunction with the disappearance of the cusp. We interpret this resonance as a virtual bound state in the limit that the two hadron scattering channels are decoupled, which correspondingly enhances the ${D^{\ast}{D}^{\ast}}$ $S$-wave amplitude upon opening. However, due to the open ${DD^{\ast}}$ channel and nonzero coupling, the pole is moved to a complex energy, manifesting as a resonance which is responsible for the cusp in the ${DD^{\ast}}$ $S$-wave amplitude.
Also examined is the behavior of the sheet II and sheet III poles as the hadron-hadron channels are decoupled. Table 6 shows the position of the sheet II and III poles in the complex ${D^{\ast}{D}^{\ast}}$ momentum plane for the different values of the coupling parameter. When the two channels are fully decoupled, the sheet III pole becomes an exact “mirror” of the sheet II pole with opposite sign of momentum. Due to this behavior and in conjunction with its distance from the physical sheet, we do not prescribe it any physical significance.
@@ V.4 Poles in other $J^{P}$
where the quoted uncertainties are statistical. We stress that this is just a single parameterisation and so no firm conclusions should be drawn. With such a small amount of constraint, higher order amplitudes can easily introduce complex poles on the physical sheet [^5] This occurs in all amplitudes attempted with a $K$-matrix with a linear polynomial in $s$.. Given that relatively few energy levels are available, we do not investigate this further. However, it is possible this resonance pole is a feature in $J^{P}=0^{-}$ ${DD^{\ast}}\{{\>\!}^{{3}\!}{P}_{0}\}$ scattering. The width is large, more than 300 MeV in physical units.
Other more distant poles can be found in some of the other partial waves. We do not consider any of these important as they are far away from where we constrain the scattering amplitudes and have a negligible effect on physical scattering.
@@ VI.1 Systematic Effects and Left-Cut Physics
The effects of left-cut physics arising from meson exchange in the scattering $t,u$-channels have not been accounted for in this calculation. They are also not currently included in the most general finite volume formalism applicable to coupled-channel scattering. The finite volume energy levels should include the necessary physics, but this has not been included in translating the lattice QCD energy levels into scattering information. At the masses used in this calculation, a $u$-channel cut due to $\pi$ exchange [^6] This is the equivalent of the “short nucleon cut” cut seen in $N\pi$ scattering ends at $a_{t}E_{cm}=0.68432$, roughly 18 MeV below threshold, and only a little below the lowest energy level extracted ($a_{t}E=0.68474(46)$ in $[000]T_{1}^{+}$ on the $L/a_{s}=16$ lattice). The virtual bound state pole we have identified would lie on this cut. The effect occurs again a similar distance below ${D^{\ast}{D}^{\ast}}$ threshold at $a_{t}E_{cm}=0.70591$. Similarly to the virtual bound state, the resonance $T_{cc}^{\prime}$ identified is located at an energy close to the opening of this cut and it is possible that including the effects of left-cut physics in this calculation could change the properties of this resonance [66]. Some modification is likely required to account for these cuts; at the very least, small imaginary parts must be introduced in the scattering amplitudes. Efforts are underway to add the necessary terms to the two-body Lüscher formalism to include these effects [67, 68], though the formalism of these modifications is not yet mature enough to handle the coupled-channel case applicable to the $T_{cc}^{\prime}$. A second alternative for the $T_{cc}$ appears to be to continue the three-body formalism a short way below ${DD^{\ast}}$ threshold [69], however this is not applicable to the resonance $T_{cc}^{\prime}$, which would require four-body dynamics. As the amplitudes are consistent with the finite volume spectrum, it is possible that the effects of the left-cut physics are not that large.
Lippmann-Schwinger and effective field theory methods applied to the ${DD^{\ast}}$ system at larger-than-physical pion masses, where the $D^{\ast}$ is stable, can be found in Refs. [66, 70, 43]. There it is shown that the presence of the left-cut below ${DD^{\ast}}$ threshold arising from one pion exchange affects both the nature of the $T_{cc}$ and its position. Similar complications can be expected below ${D^{\ast}{D}^{\ast}}$ threshold for the $T_{cc}^{\prime}$, and so extra work is needed to understand how the structures observed in this study are affected.
@@ VI.2 Comparison to Previous Work and Phenomenology
Other works have studied this channel. Considering only elastic scattering, Ref. [42] working at a light-quark mass corresponding to $m_{\pi}=280$ MeV, finds a virtual bound state 9.9 MeV below ${DD^{\ast}}$ threshold in $S$-wave. Using the HAL QCD method with almost-physical pions, Ref. [40] finds a virtual bound state much closer to threshold. While these results point toward a ground-state $T_{cc}$ that gets closer to threshold with decreasing light-quark mass, further work is needed to understand the structures in the complex plane before firm conclusions can be made. In Ref. [57], the spectrum of the $I=0$ doubly charmed sector was computed and no strong indications of a bound state or narrow resonance were present. However, the authors noted that “the situation is not as straightforward when one considers coupled-channel scattering or broad resonances”. The identification of both the $T_{cc}$ and the $T_{cc}^{\prime}$ highlights the importance of extracting enough energy levels to perform a scattering analysis in order to elucidate the pole structure in coupled-channel systems.
@@ VII Summary
In summary, we have presented the first coupled-channel ${DD^{\ast}}$, ${D^{\ast}{D}^{\ast}}$ scattering calculation from lattice QCD in $S$-wave for isospin-0. Using 109 energy levels, we determined the $J^{P}=1^{+}$ scattering amplitude by constraining other waves in the ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ channels for 12 parameterization variations that adequately describe the finite volume spectrum. In all successful variations, we observe a pole in the $S$-wave amplitudes below ${DD^{\ast}}$ threshold corresponding to a virtual bound state at approximately $\sqrt{s}\approx 3834(31)$ MeV, commonly identified as the $T_{cc}$. A pole corresponding to a resonance $T_{cc}^{\prime}$ is also found just below ${D^{\ast}{D}^{\ast}}$ threshold at $\sqrt{s}\approx 3971(35)$ MeV with a width of $\Gamma\approx 11(13)$ MeV, which predominantly couples to the kinematically closed ${D^{\ast}{D}^{\ast}}$ channel. The presence of the nearby resonance pole below ${D^{\ast}{D}^{\ast}}$ threshold enhances the ${DD^{\ast}}$ $S$-wave amplitude, resulting in clear cusp at ${D^{\ast}{D}^{\ast}}$ threshold. As ongoing experiments are currently investigating the isospin-0 $J^{P}=1^{+}$ charm sector, it should be possible to find the experimental $T_{cc}^{\prime}$ in both ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ final states.
@@ Acknowledgements.
We thank our colleagues within the Hadron Spectrum Collaboration (www.hadspec.org), particularly Jozef Dudek, Raul Briceño and Andrew Jackura for useful discussions. We would also like to thank Mikhail Mikhasenko for many insightful discussions. TW & DJW acknowledge support from a Royal Society University Research Fellowship. TW, DJW & CET acknowledge support from the U.K. Science and Technology Facilities Council (STFC) [grant numbers ST/T000694/1, ST/X000664/1]. TW also acknowledges support from Science Foundation Ireland [grant number 21/FFP-P/10186] The software codes Chroma [73], QUDA [74, 75], QUDA-MG [76], QPhiX [77], MG_PROTO [78], QOPQDP [79, 80], and Redstar [81] were used. Some software codes used in this project were developed with support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing (SciDAC) program; also acknowledged is support from the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration. This work used the Cambridge Service for Data Driven Discovery (CSD3), part of which is operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk) on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 was funded by BEIS capital funding via STFC capital grants ST/P002307/1 and ST/R002452/1 and STFC operations grant ST/R00689X/1. Other components were provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/P020259/1). This work also used the earlier DiRAC Data Analytic system at the University of Cambridge. This equipment was funded by BIS National E-infrastructure capital grant (ST/K001590/1), STFC capital grants ST/H008861/1 and ST/H00887X/1, and STFC DiRAC Operations grant ST/K00333X/1. DiRAC is part of the National E-Infrastructure. This work also used clusters at Jefferson Laboratory under the USQCD Initiative and the LQCD ARRA project. Propagators and gauge configurations used in this project were generated using DiRAC facilities, at Jefferson Lab, and on the Wilkes GPU cluster at the University of Cambridge High Performance Computing Service, provided by Dell Inc., NVIDIA and Mellanox, and part funded by STFC with industrial sponsorship from Rolls Royce and Mitsubishi Heavy Industries. Also used was an award of computer time provided by the U.S. Department of Energy INCITE program and supported in part under an ALCC award, and resources at: the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725; the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231; the Texas Advanced Computing Center (TACC) at The University of Texas at Austin; the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation Grant No. ACI-1548562; and part of the Blue Waters sustained-petascale computing project, which is supported by the National Science Foundation (awards OCI-0725070 and ACI-1238993) and the state of Illinois. Blue Waters is a joint effort of the University of Illinois at Urbana-Champaign and its National Center for Supercomputing Applications.
Data Availability
@@ Appendix A Subduction Tables
In this Appendix, we list the pseudoscalar-vector and identical vector-vector partial waves that contribute to the finite volume irreducible representations in Tables 7-10. Note that for identical vector-vector scattering, certain partial waves are not allowed by Bose symmetry (see Table 2) and these are ommitted from the tables.
@@ Appendix B Operator Tables
${D^{\ast}{D}^{\ast}}$ Partial Wave $\gamma^{(0)}_{{D^{\ast}{D}^{\ast}}\{{\>\!}^{{2S+1}\!}{\ell}_{J}\}\rightarrow{D^{\ast}{D}^{\ast}}\{{\>\!}^{{2S+1}\!}{\ell}_{J}\}}$ $\chi^{2}/N_{\mathrm{dof}}$